# Psychologically Plausible Deep Learning with Differentiable Tversky Similarity

### v1

@@ Abstract
Work in psychology has highlighted that the geometric model of similarity standard in deep learning is not psychologically plausible because its metric properties such as symmetry do not align with human perception of similarity. In contrast, Tversky (1977) proposed an axiomatic theory of similarity with psychological plausibility based on a representation of objects as sets of features, and their similarity as a function of their common and distinctive features. However, this model of similarity has not been used in deep learning before, in part because of the challenge of incorporating discrete set operations. In this paper, we develop a differentiable parameterization of Tversky’s similarity that is learnable through gradient descent, and derive basic neural network building blocks such as the Tversky projection layer, which unlike the linear projection layer can model non-linear functions such as xor. Through experiments with image recognition and language modeling neural networks, we show that the Tversky projection layer is a beneficial replacement for the linear projection layer (also known as fully connected layer), which employs geometric similarity. For instance, on the NABirds image classification task, a frozen ResNet-50 adapted with a Tversky projection layer achieves a 24.7% relative accuracy improvement over the linear layer adapter baseline. With Tversky projection layers, GPT-2’s perplexity on PTB decreases by 7.5%, and its parameter count by 34.8%. Finally, we propose a unified interpretation of both types of projection layers as computing similarities of input stimuli to learned prototypes for which we also propose a novel visualization technique highlighting the interpretability of Tversky projection layers. Our work offers a new paradigm for thinking about the similarity model implicit in modern deep learning, and designing neural networks that are interpretable under an established theory of psychological similarity.
@@ 2.2 Differentiable Tversky Similarity
Feature Set Difference: $f(A-B)$ is a measure of features present in $A$ but not present in $B$ (Equation 4). We propose an alternate form of this measure that accounts for features that are present in both $A$ and $B$, but in greater amount in $A$ (Equation 5). These two measures of set difference respectively correspond to the values ignorematch and substractmatch of the difference reduction hyperparameter of Tversky neural modules.
$f(A-B)$ $=\sum\nolimits_{k=1}^{|\Omega|}\bigl{(}a\cdot f_{k}\bigr{)}\times\mathbbm{1}\bigl{[}a\cdot f_{k}>0\land b\cdot f_{k}\leq 0\bigr{]}$ (4)
$f(A-B)$ $=\sum\nolimits_{k=1}^{|\Omega|}\bigl{(}a\cdot f_{k}-b\cdot f_{k}\bigr{)}\times\mathbbm{1}\bigl{[}a\cdot f_{k}>0\land b\cdot f_{k}>0\land a\cdot f_{k}>b\cdot f_{k}\bigr{]}$ (5)
@@ 2.4 Interpretation of Projection Layers:
Both the linear and Tversky projection layers output vectors in which each dimension is the similarity of the input to a prototype; however, they differ in the employed similarity function. Linear projection layers compute dot-product similarities. This is an application of the geometric model of smilarity, in which the features of objects and prototypes are scalar values arranged in a cartesian coordinate space and similarity is measured as the oriented length of the object vector’s projection onto the prototype vector. Dot product similarities have metric properties such as asymmetry and triangle inequality that have been empirically shown to not align with human perception, and proven to be incapable of modeling asymmetric relations. Tversky projection layers compute Tversky similarity with learnable parameters prototype vectors $\Pi$, feature vectors $\Omega$, and Tversky’s contrast model’s weights $\alpha,\beta,and\theta$. The number of features $|\Omega|$ can be varied without affecting the output dimensionality $p=|\Pi|$.
@@ 3 Experiments
Here we demonstrate the utility of our proposed differentiable Tversky similarity function in machine learning. First, we show by construction that a single Tversky projection layer can model the xor function and report empirical results on the same learning task (Section 3.1). Then, in Sections 3.2 and 3.3 we conduct empirical experiments with state-of-the-art neural networks for image recognition and language modeling in which we compare baseline neural networks with their counterparts employing Tversky projection layers. Finally, in Section 3.4, we conduct qualitative analyses employing our proposed data-domain visualization methods for projection parameters to qualitatively compare linear and Tversky projection layers and demonstrate the interpretability of Tversky neural networks.
@@ 3.1 Modeling XOR with a Single Tversky Projection Layer
As a first experiment, we construct a single Tversky projection layer that computes the xor function, which is not computable by a single linear projection due to the required non-linear decision boundary. Figure 1 shows the constructed projection, and its data and parameter vectors along with their set-centric interpretation. In empirical experiments in which we train a Tversky projection to learn xor with gradient descent under various hyperparameter conditions, we find that: Some initializations of prototypes and features lead to convergence failure; Initializing those parameters from a uniform distribution leads to higher convergence probability compared to normal and orthogonal initialization; Normalizing prototype and object vectors deteriorates convergence probability; product and substractmatch work best as values of the intersection reduction and difference reduction hyperparameters; Convergence probability doesn’t increase monotonically with the feature bank size; and finally, Tversky projection can model xor with as little as one feature. See Appendix D.
@@
Figure 1: (left) A single tversky projection layer with 11 learnable parameters that computes the xor function. The input is a 2 digit binary number encoded as $x^{i}\in\mathcal{R}^{2}$. The output $\hat{y}^{i}$ is such that $txor(x^{i})\iff s(x^{i},p^{1})>s(x^{i},p^{0})$, where $s$ is the proposed similarity function. (middle) Input vectors $x^{i}$ along with features $f^{i}$, and prototypes $p^{i}$ modeling the $xor$ function. The dot products $x^{1}\cdot f^{1}$, $x^{2}\cdot f^{0}$, $p^{1}\cdot f^{0}$ and $p^{1}\cdot f^{1}$ are positive. (right) Feature vectors define a universe in which input vectors $x^{0}=\{\}$, $x^{1}=\{f^{1}\}$, $x^{2}=\{f^{0}\}$, $x^{3}=\{\}$, and learned prototype vectors $p^{0}=\{\}$, $p^{1}=\{f^{0},f^{1}\}$, are represented as set of features. By convention, an object (input or prototype) contains a feature if its dot product with the feature vector is positive. By Tversky’s contrast model of similarity, $s(x^{3},p^{0})>s(x^{3},p^{1})$, and $s(x^{2},p^{1})>s(x^{2},p^{0})$ therefore $txor([1,1])=0$ and $txor([1,0])=1$. Similarly for other inputs, $txor([0,0])=0$ and $txor([0,1])=1$.
@@ 3.2.2 Method
ResNet-50 and TverskyResNet-50, which uses a Tversky Projection instead of the final fully connected layer, are trained with the same hyperparameters (200 epochs; batch sizes: (NABirds, 256), (MNIST, 1024); lr: 0.03; momentum 0.9; weight decay $10^{-8}$; dropout: 0.1; and gradient clipping threshold $10$). In the Pretrained backbone condition, ResNet50 is initialized with pretrained ImageNet weights, instead of random weights. In the Freeze Backbone condition, the convolution layer weights are not updated during training. Tversky prototypes and features are randomly initialized in all settings. Tversky feature intersection and differences are reduced with product and ignorematch.
@@ 3.2.3 Results
Table 1 shows trained model accuracies for each experimental condition. TverskyResNet performed better in all conditions for both datasets. In particular, when Tversky projection was used as a domain adapter with a frozen backbone, it showed $8.5\%$ and $24.7\%$ improvement over baseline. In this experiment, the Tversky neural networks have more parameters than the baseline, because their parameters also include a feature bank in addition to the prototype bank while the baseline only contains a prototype bank, corresponding to the weights of the fully connected layer. Figure 10 shows TverskyResNet-50’s accuracy on MNIST in comparison to baseline with various feature counts. TverskyResNet-50’s feature bank size is set to 224 in the NABirds experiments. While the increased parameter count of TverskyResNet-50 could contribute to the observed accuracy improvement, our results in the next sections show that the use of Tversky neural networks can simultaneously result in decreased parameter count and increased performance compared to baseline.
@@
3.3 Language Modeling with TverskyGPT-2
Table 2: Perplexity (PPL) and parameter counts (Params) of GPT2 and TverskyGPT2 trained for the language modeling task on the PTB dataset.
Model Init Tied Proto Features Params PPL $\Delta Params$ $\Delta PPL$ baseline finetuned N 163 037 184 34.85 tversky-head finetuned N 16 384 175 620 099 32.31 +7.7% -7.3% baseline finetuned Y 124 439 808 19.99 tversky-head finetuned Y 32 768 149 605 635 20.46 +20.2% +2.4% baseline scratch N 163 037 184 134.06 tversky-all-1layer scratch N 4 096 116 591 655 117.59 -28.5 % -12.3% baseline scratch Y 124 439 808 136.04 tversky-all-1layer scratch Y 8 192 81 140 007 125.86 -34.8% -7.5%
3.3.1 Task and Data
@@ 3.3.2 Method
The following TverskyGPT-2 variants were considered: tversky-head, which replaces GPT-2’s language modeling head’s stack of linear layers with a Tversky projection layer; tversky-all, which also replaces GPT-2’s intermediate attention blocks output projections with Tversky projection layers. The 2layer variant stacks two Tversky projection layers. Tversky feature sharing is employed across all Tversky projection layers. We also experiment with prototype-tying in both the baseline GPT-2 model and the Tversky models. For Tversky models, we experiment with various feature bank sizes ranging from 256 to 32,768. Finally, we experiment with initializing the models from scratch and from OpenAI’s released weights. All models were trained for 50 epochs.
@@ 3.3.3 Results
Tversky neural networks matched or surpassed baseline perplexity in all settings except when initializing the model from pretrained weights and employing prototype tying (Table 2). In that setting, Tversky prototypes and features are still randomly initialized, while the baseline fully connected layers are initialized from pretrained token embedding matrix. The perplexity gap between the best Tversky neural network and baseline was higher when training from scratch, with a 12.3% reduction in perplexity when prototypes are tied to input embeddings, and a 7.5% reduction when they are not. In the tied prototype setting, the best Tversky model also has 34.8% fewer parameters. See Appendix E for detailed results.
3.4 Qualitative Analysis of Tversky Projection Layers
Figure 2: Visualization of projection parameters using our input domain projection parameter specification method. Top: 3 examples for each class. Middle: 10 columns of the final linear projection layer of VisualMNISTNet. Bottom: 10 prototype vectors and 20 feature vectors of the Tversky Projection Layer of TverskyMNISTNet. Both models achieve 99% accuracy on the validation set. Handwritten digit features are more perceptible to humans in Tversky prototypes and features compared to the linear projection prototypes.
3.4.1 Task and Data
In this experiment we qualitatively compare the final linear projection layer of a stack of 3 linear projection layers separated by non-linear activation functions with a single Tversky projection layer in neural networks trained to perform handwritten digit recognition on the MNIST dataset using our proposed visualization method of projection layers.
3.4.2 Method
Two neural networks, VisualMNISTNet, and VisualTverskyMNISTNet (Figure 7) are trained to perform MNIST handwritten digit classification. The two neural network architectures share the same convolutional feature extraction stack yielding a 36-dimensional vector representation of the input image. They differ in how those vectors are projected onto the 10-dimensional output vectors corresponding to 10 digit classes. VisualMNISTNet employs a stack of 120, 84, and 10 unit linear projection layers following LeCun et al. (1998). However, the ReLU activation function is used instead of the logistic sigmoid. VisualTverskyMNISTNet employs a single Tversky Projection layer with 10 prototypes and 20 features. Both neural networks employ the method described in Section 2.5, facillitating our qualitative analysis of the learned prototypes and features.
3.4.3 Results
@@
Figure 3: Examples of MNIST digits sorted from low to high salience as calculated using Equation 2. This aligns with Tversky’s hypothesis that more salient stimuli exhibit more goodness of form.
Learned Tversky prototyes and features are recognizable.
They exhibit stroke patterns similar to human handwriting such as lines and curves more prominently. In contrast, the prototypes learned by the state-of-the art linear projection layer exhibit texture patterns that are hardly recognizable. This uninterpretability, even in this setting in which background textures are abscent constitutes a signifiant limitation of the prior art.
Tversky projections are explainable.
The Tversky Projection layer is based on a principled and tested theory of psychological similarity which states that humans assess the similarity between object as a linear combination of their common and distinctive features. As such, it permits explaining the similarity of one object or prototype to the other in terms of measures of their common and distinctive features as illustrated in Figure 4. This capability enables psychologically plausible understanding of neural network decisions, which is a significant improvement over state-of-the art.
Figure 4: Two MNIST digits and their learned common and distinctive features. Tversky neural networks allow explaining the similarity of objects in terms of their common and distinctive features.
Tversky Prototypes are Interpretable
Tversky’s hypothesized relation between prototypically and saliency is measurable in Tversky neural networks. The salience of objects and prototypes can be calculated at inference time using Equation 2. It is also visible in Figure 2 that prototypes are more salient than instances; they exhibit more features, and appear to combine the features of all possible instances of their class. For instance, the prototype for the handwritten 7 has both an upper-left serif and a middle horizontal stroke. The prototype for handwritten 1 and 9 have both vertical lines slanted to the left and right. The prototype for class 9 also has both an open top loop and a closed top loop.

### v2

# Tversky Neural Networks: Psychologically Plausible Deep Learning with Differentiable Tversky Similarity

@@ Abstract
Work in psychology has highlighted that the geometric model of similarity standard in deep learning is not psychologically plausible because its metric properties such as symmetry do not align with human perception of similarity. In contrast, Tversky (1977) proposed an axiomatic theory of similarity with psychological plausibility based on a representation of objects as sets of features, and their similarity as a function of their common and distinctive features. This model of similarity has not been used in deep learning before, in part because of the challenge of incorporating discrete set operations. In this paper, we develop a differentiable parameterization of Tversky’s similarity that is learnable through gradient descent, and derive basic neural network building blocks such as the Tversky projection layer, which unlike the linear projection layer can model non-linear functions such as xor. Through experiments with image recognition and language modeling neural networks, we show that the Tversky projection layer is a beneficial replacement for the linear projection layer. For instance, on the NABirds image classification task, a frozen ResNet-50 adapted with a Tversky projection layer achieves a 24.7% relative accuracy improvement over the linear layer adapter baseline. With Tversky projection layers, GPT-2’s perplexity on PTB decreases by 7.8%, and its parameter count by 34.8%. Finally, we propose a unified interpretation of both types of projection layers as computing similarities of input stimuli to learned prototypes for which we also propose a novel visualization technique highlighting the interpretability of Tversky projection layers. Our work offers a new paradigm for thinking about the similarity model implicit in modern deep learning, and designing neural networks that are interpretable under an established theory of psychological similarity.
@@ 2.2 Differentiable Tversky Similarity
Feature Set Difference: f(AB)f(A-B) is a measure of features present in $A$ but not present in $B$ (Equation 4). We propose an alternate form of this measure that accounts for features that are present in both $A$ and $B$, but in greater amount in $A$ (Equation 5). These two measures of set difference respectively correspond to the values ignorematch and substractmatch of the difference reduction hyperparameter of Tversky neural modules.
$f^{i}(A-B)$ $=\sum\nolimits_{k=1}^{|\Omega|}\bigl(a\cdot f_{k}\bigr)\times\mathbbm{1}\bigl[a\cdot f_{k}>0\land b\cdot f_{k}\leq 0\bigr]$ (4)
$f^{s}(A-B)$ $=f^{i}(A-B)+\sum\nolimits_{k=1}^{|\Omega|}\bigl(a\cdot f_{k}-b\cdot f_{k}\bigr)\times\mathbbm{1}\bigl[b\cdot f_{k}>0\land a\cdot f_{k}>b\cdot f_{k}\bigr]$ (5)
@@ 2.4 Interpretation of Projection Layers:
Both the linear and Tversky projection layers output vectors in which each dimension is the similarity of the input to a prototype; however, they differ in the employed similarity function. Linear projection layers compute dot-product similarities. This is an application of the geometric model of simlarity, in which the features of objects and prototypes are scalar values arranged in a cartesian coordinate space and similarity is measured as the oriented length of the object vector’s projection onto the prototype vector. Dot product similarities have metric properties such as asymmetry and triangle inequality that have been empirically shown to not align with human perception, and proven to be incapable of modeling asymmetric relations. Tversky projection layers compute Tversky similarity with learnable parameters prototype vectors $\Pi$, feature vectors $\Omega$, and Tversky’s contrast model’s weights $\alpha,\beta,and\theta$. The number of features $|\Omega|$ can be varied without affecting the output dimensionality $p=|\Pi|$.
@@ 3 Experiments
Here we demonstrate the utility of our proposed differentiable Tversky similarity function in machine learning. First, we show by construction that a single Tversky projection layer can model the xor function and report empirical results on the same learning task (Section 3.1). Then, in Sections 3.2 and 3.3, we conduct empirical experiments with state-of-the-art neural networks for language modeling and image recognition in which we compare baseline neural networks with their counterparts employing Tversky projection layers. Finally, in Section 3.4, we conduct qualitative analyses demonstrating the interpretability of Tversky neural networks.
@@ 3.1 Modeling XOR with a Single Tversky Projection Layer
As a first experiment, we construct a single Tversky projection layer that computes the xor function, which is not computable by a single linear projection due to the required non-linear decision boundary. Figure 1 shows the constructed projection, and its data and parameter vectors along with their set-centric interpretation. In empirical experiments in which we train a Tversky projection to learn xor with gradient descent under various hyperparameter conditions, we find that: Some initializations of prototypes and features lead to convergence failure; Initializing those parameters from a uniform distribution leads to higher convergence probability compared to normal and orthogonal initialization; Normalizing prototype and object vectors deteriorates convergence probability; product and substractmatch work best as values of the intersection reduction and difference reduction hyperparameters; Convergence probability doesn’t increase monotonically with the feature bank size; and finally, Tversky projection can model xor with as little as one feature. See Appendix C.
@@
Figure 1: (left) A single tversky projection layer with $1$ learnable parameters that computes the xor function. The input is a 2 digit binary number encoded as $x^{i}\in\mathcal{R}^{2}$. The output $\hat{y}^{i}$ is such that $txor(x^{i})\iff s(x^{i},p^{1})>s(x^{i},p^{0})$, where $s$ is the proposed similarity function. (middle) Input vectors $x^{i}$ along with features $f^{i}$, and prototypes $p^{i}$ modeling the xorxor function. The dot products $x^{1}\cdot f^{1}$, $x^{2}\cdot f^{0}$, $p^{1}\cdot f^{0}$ and $p^{1}\cdot f^{1}$ are positive. (right) Feature vectors define a universe in which input vectors $x^{0}=\{\}$, $x^{1}=\{f^{1}\}$, $x^{2}=\{f^{0}\}$, $x^{3}=\{\}$, and learned prototype vectors $p^{0}=\{\}$, $p^{1}=\{f^{0},f^{1}\}$, are represented as set of features. By convention, an object (input or prototype) contains a feature if its dot product with the feature vector is positive. By Tversky’s contrast model of similarity, $s(x^{3},p^{0})>s(x^{3},p^{1})$, and $s(x^{2},p^{1})>s(x^{2},p^{0})$ therefore $txor([1,1])=0$ and $txor([1,0])=1$. Similarly for other inputs, $txor([0,0])=0$ and $txor([0,1])=1$.
3.2 Language Modeling with TverskyGPT-2
Table 1: Perplexity (PPL) and parameter counts (Params) of GPT2 and TverskyGPT2 trained for the language modeling task on the PTB dataset. Perplexities are reported on the held out test split.
Model Init Tied Proto Features Params PPL $\Delta Params$ $\Delta PPL$ baseline finetuned N 163 037 184 30.52 tversky-head finetuned N 16 384 175 620 099 28.33 +7.7% -7.2% baseline finetuned Y 124 439 808 18.31 tversky-head finetuned Y 32 768 149 605 635 18.62 +20.2% +1.7% baseline scratch N 163 037 184 111.79 tversky-all-1layer scratch N 4 096 116 591 655 98.22 -28.5 % -12.1 % baseline scratch Y 124 439 808 112.81 tversky-all-1layer scratch Y 8 192 81 140 007 103.99 -34.8% -7.8 %
3.2.1 Task and Data
3.2.2 Method
The following TverskyGPT-2 variants were considered: tversky-head, which replaces GPT-2’s language modeling head’s stack of linear layers with a Tversky projection layer, and tversky-all, which also replaces GPT-2’s intermediate attention blocks’ output projections with Tversky projections. Tversky feature sharing is employed across all Tversky projection layers. For all models, we experiment with prototype tying (tied vs not tied) and initialization (random vs OpenAI’s released weights) while training for 50 epochs on PTB’s training set, and validating on PTB validation set. The baseline models, and best Tversky models are evaluated on PTB’s held-out test set.
3.2.3 Results
Tversky language models matched or surpassed baseline perplexity in all settings except the one in which pre-trained weights and prototype tying are employed (Table 1). In that setting, Tversky prototypes and features are still randomly initialized, while the baseline fully connected layers are initialized from the pretrained token embedding matrix. The perplexity gap between the best Tversky neural network and baseline was higher when training from scratch, with a 7.8% reduction in perplexity when prototypes are tied to input embeddings, and a 12.1% reduction when they are not. In the tied prototype setting, the best Tversky model also has 34.8% fewer parameters. See Appendix D for validation results, and Table 1 for test results.
@@ 3.3.2 Method
ResNet-50 and TverskyResNet-50, which uses a Tversky Projection instead of the final fully connected layer, are trained with the same hyperparameters (200 epochs; batch sizes: (NABirds, 256), (MNIST, 1024); lr: 0.03; momentum 0.9; weight decay $10^{-8}$; dropout: 0.1; and gradient clipping threshold $10$). In the Pretrained backbone condition, ResNet50 is initialized with pretrained ImageNet weights. In the Freeze Backbone condition, the convolution layer weights are not updated during training. Tversky prototypes and features are randomly initialized in all settings. Tversky feature intersection and difference reduction hyperparameters (product and ignorematch) were informed by our results in Section 3.2. We chose Tversky feature bank sizes 20 and 224 for MNIST and NABirds respectively. Unlike in experiment 3.2, which employs a strict train/validation/test protocol, no extensive hyperparameter tuning, and validation was employed in this experiment because the employed datasets, despite having sizeable test sets, do not have official distinct validation sets. Nonetheless the following results demonstrate the existence of settings under which Tversky vision models match or surpass baseline accuracy on official train/test splits.
@@ 3.3.3 Results
Classification accuracies reported in table 2 show that Tversky vision models can match or surpass baseline accuracy under the specified experimental conditions. In particular, when Tversky projection was used as a domain adapter with a frozen backbone, $8.5\%$ and $24.7\%$ improvement over baseline was observed. In this experiment, the Tversky neural networks have more parameters than the baseline, because their parameters also include a feature bank in addition to the prototype bank while the baseline only contains a prototype bank, corresponding to the weights of the fully connected layer. While the increased parameter count of TverskyResNet-50 could contribute to the observed accuracy improvement, our results in Section 3.2 show that the use of Tversky neural networks can simultaneously result in decreased parameter count and increased performance compared to baseline.
@@
3.4 Qualitative Analysis of Tversky Neural Networks
Figure 2: Visualization of prototypes using our input domain projection parameter specification method. Top: 3 examples for each class. Bottom: 10 columns of the final linear projection layer of VisualMNISTNet, and 10 prototype vectors of the Tversky Projection Layer of TverskyMNISTNet. Both models achieve 99% accuracy on the validation set. Handwritten digit features are more perceptible to humans in Tversky prototypes and features compared to the linear projection prototypes.
Here we present results showing the interpretability of Tversky neural networks. We discuss the explainability of learned prototypes, the quantification of salience, and the algebraic specification of semantic fields enabled by Tversky neural networks.
@@ 3.4.1 Method
Prototype Visualization: To visually compare the prototypes learned by Tversky Neural Networks and classical neural networks, we employ the prototype visualization method described in Section 2.5 and illustrated in Appendix B. This method is used to display the handwritten digit prototypes learned by VisualMNISTNet and VisualTverskyMNISTNet, two neural networks inspired by LeNet-5 (LeCun et al., 1998). Further details are provided in Appendix F.
Salience: To assess whether salience, as computed using Equation 2, is intelligible to humans, we rank MNIST examples by their salience score.
Semantic Fields: Tversky Neural Networks’ repesentation of objects as sets of features allows the algebraic specification of semantic fields. A semantic field is a subset of features resulting from set expressions that are interpretable by humans. For instance, $A\cap B-C$ describes the semantic field capturing the common features of $A$ and $B$ that are distinct from $C$. To visualize a semantic field $F$, we rank examples $x_{i}$ by their score in that field: $s=\sum_{f_{k}\in F}{f_{k}\cdot x_{i}}$. This enables the visualization of semantics captured by TverskyMNISTNet’s representation of images and TverskyGPT2 representation of tokens.
@@ 3.4.2 Results
Prototypes: Figure 2 shows that Tversky prototypes exhibit stroke patterns similar to human handwriting, such as lines and curves, more clearly than those learned by linear projection layers, which exhibit texture patterns that are difficult to interpret. This lack of interpretability, even in this simple domain without background textures, represents a significant limitation of prior approaches. Tversky prototypes are also more salient than individual instances: they exhibit more features and appear to combine the features of all possible instances of their class. For example, the prototype for the handwritten “7” has both an upper-left serif and a middle horizontal stroke. The prototypes for the digit “1” and “9” include vertical lines slanted to the right and to the left. The prototype for class “9” combines both open and closed top loops. Figure 15 shows the evolution of learned prototypes over training epochs for both VisualMNISTNet and TverskyVisualMNISTNet. We observed that $\alpha>\beta$ in trained Tversky neural networks (see Appendix G), indicating that they weigh the distinctive features of individual instances more heavily than those of prototypes, which is in line with Tversky’s theory.
@@
Salience: Figure 3 shows MNIST examples ranked from low to high salience according to Equation 2. This ranking aligns with Tversky’s empirical psychological observations that humans perceive stimuli exhibiting more goodness of form as more salient.
Semantic Fields Table 3 shows examples of semantic fields defined with TverskyGPT2 tokens, and top scoring tokens in those semantic fields. These results and additional results in Appendix E show that interpretable semantic fields such as those capturing the concepts of adjectives, comparatives, superlatives, verb forms, or specific word senses can be specified algebraically with Tversky Language models. Similar results can be achieved with Tversky vision models. Figure 16 shows two MNIST digits, and examples that illustrate their common and distinctive features.
Table 3: Semantic concepts, their algebraic specification using set operations on TverskyGPT2 tokens represented as sets, and top tokens ranked by semantic score in the resulting semantic field. Tversky neural networks enable the algebraic specification of interpretable semantic concepts.
Concept Semantic Expression Top-scoring tokens
Adjective $bad\cap good-worse-better-worst-best$ lousy, evil, crappy, poor, shitty, nice, terrible, horrible, decent, mediocre, valid, great, excellent
Comparative $worse\cap better-bad-good-worst-best$ sharper, nicer, clearer, smarter, wiser, smoother, safer, hotter, happier, tighter, louder, preferable
Superlative $worst\cap best-worse-better-bad-good$ happiest, safest, quickest, deadliest, finest, hottest, busiest, toughest, darkest, fastest, brightest, holiest
Industrial Plant: plantaceaevegetationvegetableplantingflowerherbvinecroptreemushroomplantationleavesplant-aceae-vegetation-vegetable-planting-flower-herb-vine-crop-tree-mushroom-plantation-leaves facility, stall, brate, factory, Laboratories, Shed, implant, Berm, Sew, Manufacturing, reactor, Diesel, refinery, strain, distribut, planting, Industrial, microbial, actory, reactors, Unit, Install, depot, spill, Cutter, Indust, Bot, Nuclear, ineries
Figure 3: MNIST digits sorted from low to high salience (Equation 2). Like humans, the underlying Tversky Neural Network perceives stimuli exhibiting more goodness of form as more salient.